### v2

# MeanVC: Lightweight and Streaming Zero-Shot Voice Conversion via Mean Flows

Abstract
Zero-shot voice conversion (VC) aims to transfer timbre from a source speaker to any unseen target speaker while preserving linguistic content. Growing application scenarios demand models with streaming inference capabilities. This has created a pressing need for models that are simultaneously fast, lightweight, and high-fidelity. However, existing streaming methods typically rely on either autoregressive (AR) or non-autoregressive (NAR) frameworks, which either require large parameter sizes to achieve strong performance or struggle to generalize to unseen speakers. In this study, we propose MeanVC, a lightweight and streaming zero-shot VC approach. MeanVC introduces a diffusion transformer with a chunk-wise autoregressive denoising strategy, combining the strengths of both AR and NAR paradigms for efficient streaming processing. By introducing mean flows, MeanVC regresses the average velocity field during training, enabling zero-shot VC with superior speech quality and speaker similarity in a single sampling step by directly mapping from the start to the endpoint of the flow trajectory. Additionally, we incorporate diffusion adversarial post-training to mitigate over-smoothing and further enhance speech quality. Experimental results demonstrate that MeanVC significantly outperforms existing zero-shot streaming VC systems, achieving superior conversion quality with higher efficiency and significantly fewer parameters. Audio demos and code are publicly available at https://aslp-lab.github.io/MeanVC.
Index Terms—  streaming voice conversion, zero-shot, mean flows, adversarial post-training
1 Introduction
Zero-shot voice conversion (VC) aims to transfer the timbre from the source speaker to any unseen speaker while keeping the linguistic content unchanged. This capability has profound implications for a wide array of applications, including movie dubbing [bgmvc, expressive-vc], privacy protection [voiceprivacy2022], personalized pronunciation correction, and the creation of expressive virtual avatars. Benefiting from modern deep learning based approaches, recent voice conversion models [DBLP:conf/icassp/HussainNHLG23, DBLP:conf/icassp/LiLL23, DBLP:conf/icassp/SEF-VC, DBLP:conf/icassp/LuoD24, DBLP:conf/icassp/KimKCNMC25, DBLP:conf/icassp/ChoiP25] have achieved remarkable performance in terms of naturalness and speaker similarity. However, as the demand for real-time processing in various application scenarios increases, the computational cost of high-performing models becomes a significant bottleneck for practical deployment. This has created a pressing need for streaming models that are simultaneously fast (low-latency), lightweight (low computational overhead), and high-fidelity (high audio quality and speaker similarity).
Recent streaming zero-shot VC studies have explored two main architectural paradigms: autoregressive (AR) and non-autoregressive (NAR) frameworks. A representative AR-based streaming VC system is StreamVoice [DBLP:conf/acl/WangCW0W24], which employs a fully causal, context-aware language model. It alternates between semantic bottleneck features and acoustic codec tokens at each AR step, supported by an auxiliary acoustic predictor, to enable streaming zero-shot VC without future look-ahead. Building on this design, StreamVoice+ [DBLP:journals/spl/WangCWXW24] introduces a learnable semantic encoder and a connector, while adapting the backbone via LoRA [DBLP:conf/iclr/HuSWALWWC22]. It further improves quality through a residual-bottleneck connector, forming an end-to-end streaming pipeline. Despite their high-fidelity output, these models fundamentally compromise on the other two critical requirements. Their sequential decoding process inherently leads to high latency, failing the ’fast’ criterion. Moreover, their massive model sizes (101M for StreamVoice and 153M for StreamVoice+) make them far from ’lightweight,’ resulting in prohibitive computational costs that render real-time inference on standard CPUs impractical.
To mitigate the inherent latency of AR models, NAR frameworks have been explored for their potential in parallel generation. DualVC2 [DBLP:conf/icassp/NingJ0WY0B24], for instance, presents a lightweight NAR model based on the conformer architecture, which accelerates inference by incorporating dynamic chunk training. While it achieves faster speeds, its reliance on conventional architectures limits its generalization capability to unseen speakers in zero-shot scenarios. Another NAR approach adapts powerful diffusion models for zero-shot VC. Seed-VC [DBLP:journals/corr/seedvc] employs a diffusion transformer (DiT) [DBLP:conf/iccv/PeeblesX23] architecture within a sliding-window pipeline. Although this mechanism enables streaming VC, the fixed-size windows can disrupt long-term contextual dependencies and speaker characteristics. Thus, while NAR frameworks successfully address the ’fast’ requirement, they often do so by sacrificing ’high-fidelity,’ either through poor generalization in simpler models or context fragmentation in more complex ones. This leaves a clear gap in the field: a framework that can unite the speed of NAR models with the quality and consistency of their AR counterparts.
Fig. 1: Overall architecture of our proposed MeanVC.
To address these challenges, we propose MeanVC, a lightweight framework for streaming zero-shot VC that leverages the advantages of both AR and NAR paradigms while addressing their respective limitations. Specifically, MeanVC employs a chunk-wise autoregressive denoising strategy, which processes audio in individual chunks while explicitly conditioning each chunk on the previous one, thereby preserving long-term speaker consistency. To overcome the latency of iterative diffusion samplers, we introduce mean flows [DBLP:journals/corr/MeanFlows], a powerful method that enables the diffusion model to generate high-quality results with 1-NFE (number of function evaluations). We integrate mean flows into the DiT-based decoder. By training the model to regress the average velocity field of the probability flow, MeanVC can synthesize a high-quality spectrogram using 1-NFE. Finally, we incorporate a diffusion adversarial post-training (DAPT) stage to mitigate over-smoothing and further enhance speech quality. With merely 14M parameters, MeanVC achieves superior audio quality and efficiency. Both subjective and objective evaluations demonstrate that MeanVC significantly outperforms existing streaming VC systems. Audio demos and code are publicly available at https://aslp-lab.github.io/MeanVC.
2 MEANVC
2.1 Overview
As illustrated in Fig. 1, MeanVC is built on the widely used recognition-synthesis framework [DBLP:journals/taslp/ZhangLD20], consisting of a streaming automatic speech recognition (ASR) module, a speaker encoder, a timbre encoder, a DiT decoder, and a vocoder. First, the pre-trained streaming ASR model extracts bottleneck features (BNFs) from the source waveform. Subsequently, these BNFs are fed into the timbre encoder, which fuses them with fine-grained timbre information extracted from a reference mel-spectrogram to produce timbre BNFs. Meanwhile, speaker embeddings are extracted from a reference waveform via a pre-trained speaker encoder. The DiT decoder generates the converted mel-spectrogram conditioned on both the speaker embeddings and the timbre BNFs. During generation, the source mel-spectrogram is concatenated at the front to serve as a cache, enabling chunk-wise autoregressive denoising and producing the converted spectrogram with the target speaker’s timbre.
2.2 Chunk-wise Autoregressive Denoising
To enable streaming conversion, we adopt a chunk-wise autoregressive approach that divides the BNFs into smaller chunks and applies a chunk-wise causal mask inspired by MoonCast [DBLP:journals/corr/MoonCast]. This mask strategy allows each chunk to access the historical context from previously generated chunks, ensuring consistency across chunks. Furthermore, we build on conditional flow matching (CFM) to learn optimal transport paths between the two distributions. Given data $x\sim p_{\text{data}}(x)$, prior $\epsilon\sim p_{\text{prior}}(\epsilon)$, the optimal transport flow path can be constructed as: $z_{t}=(1-t)x+t\epsilon$, and the conditional velocity is thus given by $v_{t}=dz_{t}/dt=\epsilon-x$. During training, the objective is to learn a neural network $f_{\theta}$ that minimizes the following CFM loss: $\mathcal{L}_{\text{CFM}}(\theta)=\mathbb{E}_{t,x,\epsilon}\left\|f_{\theta}\left(t,z_{t}\right)-v_{t}\right\|^{2}$.
Our DiT decoder is conditioned on timbre BNFs and speaker embeddings to generate the mel-spectrogram from random Gaussian noise. Specifically, we repeat the speaker embeddings to match the length of the timbre BNFs and concatenate them to form the conditioning inputs for the model. We select chunk $i$ to be generated, along with all preceding chunks, where $i$ is an integer index and $i\in[0,N]$ with $N$ being the total number of chunks. Let $M_{i}$ denote the mel-spectrogram of chunk $i$ and $C_{i}$ denote its corresponding conditioning input. The mel-spectrograms of the preceding chunks are collectively denoted as $M_{<i}$. The flow-matching approach involves a forward process to add noise to the data and a backward process to remove the noise in reverse. During the training phase, the forward process takes as input the noisy chunk $Z_{i}(t)=(1-t)M_{i}+t\xi$, where the clean mel-spectrogram $M_{i}$ is mixed with Gaussian noise $\xi\sim\mathcal{N}(0,1)$ at timestep $t\in[0,1]$. The flow-matching model $f_{\theta}$ parameterized by $\theta$, learns the velocity field mapping $f_{\theta}(t,Z_{i}(t))=dZ_{i}(t)/{dt}=\xi-M_{i}$ conditioned on $C_{i}$.
Fig. 2: Chunk-wise causal mask in the DiT decoder. In this example, each noisy mel-spectrogram chunk $Z_{i}$ can attend to up to 3 preceding clean mel-spectrogram chunks $M_{j}$ (where $j\in[i-3,i-1]$) and itself. The green cells indicate allowed attention, while the white cells indicate restricted attention.
The mel-spectrograms of preceding chunks, denoted as $M_{<i}$, are used as prompts for in-context learning. We apply a chunk-wise causal mask, as illustrated in Fig. 2. We denote the clean mel-spectrogram chunk as $M_{i}$ and the noisy mel-spectrogram chunk used for generation as $Z_{i}$. During training, all clean chunks $M_{i}$ are concatenated before all noisy chunks $Z_{i}$ to form a complete sequence of $2N$ chunks. The chunk-wise attention mask is defined as follows: 1) Full attention is allowed within each chunk. 2) For clean chunks where $i\in[0,N)$, each chunk $M_{i}$ can only attend to itself. 3) For noisy chunks where $i\in[N,2N)$, each chunk $Z_{i}$ can attend to preceding clean chunks $M_{j}$ where $j\in[\max(0,i-K),i-1]$ as well as to itself, where $K$ represents the maximum number of preceding clean chunks accessible for conditioning. We find that when using extremely short chunks (e.g., 160 ms), allowing access to too many historical chunks causes the model to over-rely on historical context rather than the current chunk $Z_{i}$, ultimately degrading performance.
2.3 Mean Flows for Generative Modeling
CFM typically requires multiple sampling steps to solve the ordinary differential equation (ODE), and the number of function evaluations directly impacts inference efficiency. With just 1-NFE, the quality of generated spectrograms degrades catastrophically. To address this limitation, we adopt the approach proposed in mean flows [DBLP:journals/corr/MeanFlows], which regresses the average velocity field during training to enable high-quality synthesis with only 1-NFE. Specifically, given a time interval $[r,t]$, the displacement on the ODE trajectory within it is $\int_{r}^{t}v(z_{\tau},\tau)d\tau$. The average velocity is then defined as: $u(z_{t},r,t)\triangleq\frac{1}{t-r}\int_{r}^{t}v(z_{\tau},\tau)d\tau$. By differentiating both sides with respect to $t$ and subsequently rearranging the resulting terms, the relationship between $u$ and $t$ can be derived, yielding the mean flows identity:
$u(z_{t},r,t)=v(z_{t},t)-(t-r)\frac{d}{dt}u(z_{t},r,t).$ (1)
By utilizing the Jacobian-vector product (JVP), the total derivative can be expressed as: $\frac{d}{dt}u(z_{t},r,t)=v(z_{t},t)\partial_{z}u+\partial_{t}u$. The velocity $v(z_{t},t)$ is the marginal velocity in flow matching. We follow [DBLP:conf/iclr/LipmanCBNL23] to replace it with the conditional velocity $v_{t}=\epsilon-x$. With this, the target field for training is defined as $u_{\text{tgt}}=v_{t}-(t-r)(v_{t}\partial_{z}u_{\theta}+\partial_{t}u_{\theta})$. We then encourage $f_{\theta}$ to satisfy this field by minimizing the mean flows objective:
$\mathcal{L}_{\text{MF}}(\theta)=\mathbb{E}_{t,r,x,\epsilon}\left\|f_{\theta}(z_{t},r,t)-\text{sg}(u_{\text{tgt}})\right\|^{2},$ (2)
where $\text{sg}(\cdot)$ denotes the stop-gradient operation. Note, when $t=r$, the mean flows objective reduces to the standard flow matching objective.
During sampling, the time integral in CFM can be substituted with the average velocity, resulting in:
$z_{r}=z_{t}-(t-r)u(z_{t},r,t).$ (3)
For 1-NFE sampling, we have $z_{0}=z_{1}-f_{\theta}(z_{1},0,1)$, where $z_{1}=\epsilon\sim p_{\text{prior}}(\epsilon)$.
2.4 Diffusion Adversarial Post-Training
Mel-spectrograms generated by flow matching models often suffer from over-smoothing artifacts. While conventional multi-length discriminators [DBLP:journals/corr/HiFiSinger] are employed, we find that high-frequency components in generated mel-spectrograms still exhibit artifacts such as bright lines. Inspired by [DBLP:journals/corr/APT], we leverage the generator architecture as the discriminator backbone to address this issue. Specifically, we employ the DiT model as generator $G$, which generates spectrograms with 1-NFE. The DiT weights are used to initialize the backbone of the discriminator $D$. After training, the discriminator outputs a logit that distinguishes between real samples $x$ and generated samples $\hat{x}$. To achieve this, we modify the DiT architecture by introducing cross-attention-only Transformer blocks at the second and fourth layers of the 4-layer DiT backbone. These blocks use a learnable vector as the query and perform cross-attention with the backbone’s latent representations as keys and values. The outputs are concatenated channel-wise to form a global feature vector, which is then projected to produce a scalar logit output.
The training objectives of DAPT are defined as follows:
$\mathcal{L}_{\text{Adv}}(G)=\mathbb{E}_{c,\epsilon}\left\|D(G(\epsilon,c),c)-1\right\|^{2},$ (4)
$\mathcal{L}_{\text{Adv}}(D)=\mathbb{E}_{c,x}\left\|D(x,c)-1\right\|^{2}+\mathbb{E}_{c,\epsilon}\left\|D(G(\epsilon,c),c)\right\|^{2},$ (5)
where $c$ denotes the conditioning input (timbre BNFs and speaker embeddings). The generator $G$ learns to produce realistic samples through 1-NFE generation, while the discriminator $D$ learns to distinguish between real and generated samples.
3 EXPERIMENTS
3.1 Experiment Setup
Dataset. We train MeanVC on the open-source Emilia [DBLP:conf/slt/Emilia] corpus. Specifically, we use DNSMOS [DBLP:conf/icassp/Dnsmos] to filter out audio with DNSMOS scores below 3.4, retaining 10,000 hours of mandarin data. All audio files are resampled to 16 kHz for VC training. For zero-shot evaluation, we use the Mandarin subset of the Seed-TTS test set [DBLP:journals/corr/Seed-tts], comprising 2,018 source–target testing pairs. For known speakers evaluation, we fine-tune MeanVC on Aishell3 [DBLP:conf/interspeech/AISHELL-3], selecting four speakers as targets for evaluation. We use 100 clean and 100 noisy clips as source recordings. To extract semantic features, we incorporate the streaming ASR model Fast-U2++ [DBLP:conf/icassp/Fast-U2++], implemented by WeNet [DBLP:conf/interspeech/WeNet] and trained on WenetSpeech [DBLP:conf/icassp/WENETSPEECH].
Implementation Details. MeanVC consists of 14M parameters. The DiT decoder employs four DiT blocks, each with a hidden size of 512 and two attention heads. The timbre encoder incorporates two cross-attention modules, with a hidden size of 256 and four attention heads. Fast-U2++ utilizes a 160 ms chunk size for streaming inference and compresses 16 kHz waveforms into semantic features with a 40 ms frame length. Speaker embeddings are extracted using ECAPA-TDNN [DBLP:conf/interspeech/ECAPA-TDNN]. We employ Vocos [DBLP:conf/iclr/Vocos] as the vocoder to convert mel-spectrograms to high-fidelity 16 kHz speech waveforms.
Table 1: Zero-shot performance (unseen speakers). The RTF is calculated only for the VC module. The Latency represents the delay of a single chunk (160 ms) within the entire pipeline. The best-performing metrics are indicated in bold, while the second-best are underlined.
Method Quality Similarity Efficiency
[…]
GT $4±\pm0$.05 3.79 1.36 - - - - -
[…]
Evaluation Metrics. For subjective evaluation, we employ naturalness mean opinion score (NMOS) and speaker similarity mean opinion score (SMOS) as subjective metrics, calculated with 95% confidence intervals. We randomly select 100 testing pairs and recruit 15 listeners to conduct the assessments. For objective evaluation, we measure speech intelligibility using character error rate (CER) computed by an ASR model [^1] https://huggingface.co/funasr/paraformer-zh. Speaker similarity (SSIM) is computed using a WavLM-finetuned speaker verification model [^2] https://github.com/BytedanceSpeech/seed-tts-eval to measure the similarity between converted and target speaker speech. Additionally, we employ DNSMOS [^3] https://github.com/microsoft/DNS-Challenge/tree/master/DNSMOS to assess speech quality. For streaming performance evaluation, we use real-time factor (RTF) and latency metrics. RTF represents the ratio of model inference time to input speech duration, indicating computational efficiency. We evaluate both RTF and latency of all VC models on a single-core AMD EPYC 7542 CPU with single-threaded execution.
Baseline Systems. To validate the performance of MeanVC, we compare MeanVC against the following baseline systems: StreamVoice [DBLP:conf/acl/WangCW0W24], a language model-based system that employs fully causal context-awareness. It alternately processes semantic and acoustic features to achieve streaming zero-shot voice conversion; Seed-VC [DBLP:journals/corr/seedvc], a NAR system based on diffusion models. It employs an external timbre shifter to mitigate timbre leakage and achieves streaming zero-shot voice conversion through a sliding-window pipeline; DualVC2 [DBLP:conf/icassp/NingJ0WY0B24], a lightweight NAR system based on the conformer architecture. It uses classic non-causal convolutions and dynamic masking convolutions to achieve streaming any-to-many voice conversion.
3.2 Evaluation on Zero-shot VC
To evaluate zero-shot VC performance, we benchmark MeanVC against Seed-VC and StreamVoice. As shown in Table 1, MeanVC outperforms all baseline systems on both subjective SMOS and NMOS, and achieves the best performance on objective metrics CER and SSIM, highlighting its effectiveness in zero-shot VC. However, MeanVC lags behind Seed-VC in DNSMOS, which we attribute to its relatively smaller parameter size compared to Seed-VC.
Our VC module consists of only 14M parameters, achieving an RTF of 0.136 on a single-core, single-threaded setup of an AMD EPYC 7542 CPU, significantly lower than baseline systems. To meet real-time requirements, RTF must remain below 1.0. Our complete pipeline includes the ASR encoder with an RTF of 0.120 and the vocoder with an RTF of 0.066, achieving an overall RTF of 0.322 and satisfying real-time constraints. With a 160 ms chunk size, the model’s inference latency is 51.52 ms. The total pipeline latency, including the chunk duration, is 211.52 ms (160 ms + 51.52 ms), significantly outperforming baseline systems. Notably, all efficiency measurements for the experiments were obtained without optimizations such as quantization.
Table 2: Intra-dataset performance (seen speakers)
Method Clean Noise
[…]
GT 3.64 0.37 - 2.86 2.92 -
DualVC2 3.63 3.84 0.659 3.47 16.28 0.562
MeanVC 3.69 3.33 0.681 3.56 12.84 0.633
[…]
3.3 Evaluation on Intra-dataset
To gain deeper insights into MeanVC’s performance, we explore it on known speakers by fine-tuning it on Aishell3, selecting four speakers as targets for evaluation. We select DualVC2 for comparison, an any-to-many VC model. Specifically, we use a configuration with parameter size comparable to MeanVC and train DualVC2 from scratch on both Emilia (using the same filtered training data as MeanVC) and Aishell3. As shown in Table 2, MeanVC outperforms DualVC2 across all metrics even without fine-tuning, with its advantages becoming particularly evident under noisy input conditions. This demonstrates MeanVC’s superior robustness and strong conversion capabilities. With available utterances of target speakers, fine-tuning significantly improves performance. This indicates our system can be easily applied to various scenarios with or without the utterances of target speakers.
3.4 Ablation Study
To validate the contribution of MeanVC’s key components and analyze the impact of key hyperparameters, we conducted a series of ablation studies, with results detailed in Table 3. First, we assessed the model’s core modules. Ablating the DAPT stage resulted in a tangible degradation across all metrics, confirming its crucial role in mitigating over-smoothing and enhancing speech quality. Similarly, disabling the clean mel-spectrogram cache led to a significant performance drop, particularly in CER, which underscores the mechanism’s importance in reducing the train-inference mismatch to improve speech intelligibility. Next, we investigated the critical trade-off between latency and performance by varying the streaming chunk size from our 160 ms default. Reducing the chunk size to 80 ms halves the processing latency, offering a faster response. However, this came at the cost of degraded performance, especially in CER and DNSMOS, likely due to insufficient contextual information for robust modeling. Conversely, increasing the chunk size to 200 ms provided more acoustic context and yielded substantial improvements across all metrics, but at the expense of a 25% increase in overall latency. These findings explicitly demonstrate the performance-latency trade-off inherent in the streaming pipeline and validate our choice of 160 ms as a well-balanced configuration that achieves strong performance without incurring excessive delay.
Table 3: Results of ablation studies. The baseline MeanVC uses a chunk size of 160ms.
[…]
MeanVC 3.76 5.01 0.687
     w/o DAPT 3.68 5.86 0.673
     w/o clean chunks 3.71 5.97 0.677
     w/ chunk size (80ms) 3.56 9.97 0.619
     w/ chunk size (200ms) 3.83 4.42 0.700
4 CONCLUSIONS
In this study, we introduce MeanVC, a lightweight and efficient framework for streaming zero-shot voice conversion. By integrating the advantages of both autoregressive and non-autoregressive approaches, MeanVC employs a chunk-wise autoregressive denoising strategy, leverages mean flows for efficient single-step spectrogram synthesis, and incorporates diffusion adversarial post-training to enhance speech naturalness. Despite its compact architecture of only 14M parameters, MeanVC achieves superior performance in both audio quality and computational efficiency. Our extensive evaluations demonstrate that MeanVC outperforms existing streaming VC systems.